# Parametric mapping of the efficiency–instability relation in plasma-wakefield accelerators

### v1

@@ I Introduction
To be viable for colliders, the plasma-wakefield acceleration (PWFA) scheme must ensure the accelerated electron bunch is stable enough transversely to achieve high luminosity in collisions. In the blowout regime [10], where a driving electron bunch with a larger density than the background plasma expels all the plasma electrons in its path and forms a focusing and accelerating ion column, transverse stability is limited by the coupling between the beam and the electrons traveling in the sheath of the blowout. High driver-to-trailing-bunch power-transfer efficiency is also required in plasma-accelerator facilities [11] for them to operate cost-effectively. However, increasing the driver-to-trailing-bunch power-transfer efficiency gives rise to larger growth rates of transverse instabilities. Such transverse instabilities lead to emittance growth, oscillation-amplitude growth, and, if not mitigated, significant charge loss in PWFAs.
The first type of transverse instability for the blowout regime in PWFAs was discussed by Whittum et al. [12] and found to have a rapid growth rate. The instability is seeded by transverse asymmetries such as tilts in the beam. The offset beam electrons perturb the path of the plasma electrons traveling in the sheath of the plasma wake. The beam oscillates about the axis due to the focusing forces from the exposed ions. The perturbation of the plasma wake and the oscillation of the beam slices are coupled and behave as harmonic oscillators. Each longitudinal slice drives the oscillation of the slice positioned right behind it, leading to a resonance behavior whose amplitude increases exponentially towards the tail of the beam. If not mitigated, this growth continues exponentially with the propagation distance. The instability was originally discussed for single beams in plasma (i.e., the drive beam), in which case it is known as the hose instability due to its firehose-like behavior. This instability was later found to be mitigated by the inherent energy spread developed in the driver during acceleration [13], as well as the non-linear focusing fields present when the transverse beam size is large [14]. However, these mitigation techniques are not sufficient for the trailing bunch, as future collider applications typically require sub-percent energy spread [11, 15], and the trailing bunch must be matched to small bunch sizes to preserve emittance [16, 8]. When discussing the trailing bunch, this instability is often referred to as the beam-breakup instability and is conceptually similar to that discussed in conventional RF accelerators since the 1960s [17].
@@ II Mapping the parameter space with only three variables
We aim to scan the entire parameter space that affects the strength of the BBU instability. The parameters chosen can be scaled to any plasma density. We turn off ion motion and suppress the energy spread by only working with optimally loaded wakefields and assuming a cold plasma. We transversely offset the trailing bunch by 0.01/$k_{p}$, sufficiently small to perturb the shape of the plasma wake only weakly. The plasma wake has a thin electron sheath [22], such that at the longitudinal position where the blowout is at its largest, the only contribution of the driver is to determine the maximum radius of the blowout $R_{b}$, which determines the possible accelerating fields $E_{z}$ that can be loaded by the trailing bunch. The second parameter to consider is the strength of the accelerating field, which is determined by the longitudinal position of the head of the trailing bunch in the plasma wake. The third parameter is the bunch length of the accelerated bunch. When we scan the bunch length, the charge of the trailing bunch is not a free parameter, as there is only one current profile that can load a perfectly uniform field at a given accelerating gradient and wake radius. The parameters that we include in our grid are, therefore: (1) the normalized wake radius $R_{b}k_{p}$; (2) the normalized accelerating field $E_{z}/E_{0}$; and (3) the bunch length of the trailing bunch.
[…]
Figure 2: Grid schematic illustrating the parameter space (selected points only shown). In one dimension, we scan over the blowout wake size $R_{b}k_{p}$. The driver changes only in current along the $R_{b}k_{p}$ direction. In the other dimension, we scan over the normalized accelerating field that the trailing bunch is sampling $E_{z}/E_{0}$. Along the $E_{z}/E_{0}$ direction, the driver is identical, but the longitudinal offset of the trailing bunch changes and the beam current is adjusted to beam load optimally. Each simulation also provides a scan over the bunch length. The top left inset in the grey region shows the full range of achievable normalized accelerating fields for each value of $R_{b}k_{p}$ in green.
@@ IV Results
As the value of $E_{z}/E_{0}$ increases and approaches the optimal value, the value of $\eta_{t}/\eta_{t,\mathrm{predicted}}$ decreases at a rate that is roughly constant in log-log space. However, as the value of $E_{z}/E_{0}$ passes the optimal value and begins to increase again, the value of $\eta_{t}$ becomes larger at an increasing rate. In particular, $\eta_{t}/\eta_{t,\mathrm{predicted}}$ for the cases with fields larger than the optimal, increases exponentially when the power-transfer efficiency $\eta_{p}$ tends towards 1. This is shown in Fig. 4 for the normalized maximum blowout radii of $R_{b}k_{p}=0.99$ and $R_{b}k_{p}=2.55$ and power-transfer efficiencies of 5%, 20% and 50%. Fitting to the form
@@
gives very good agreement with results for values of $E_{z}/E_{0}$ smaller than the optimal, which is obtained when $\eta_{t}/\eta_{t,\mathrm{predicted}}$ is minimized.
@@
This equation indicates the optimal accelerating field for minimizing transverse instability, and works for efficiencies up to $\sim$60% [^1] We exclude the two highest $R_{b}k_{p}$ values to improve the fit.. Equation 9 predicts the simulation minima to within 10% at low efficiency and 14% at high efficiency.
@@ V Amplitude growth
The oscillation-amplitude growth of the tail of the trailing bunch in each case is evaluated against the normalized quantity $\mu\eta_{t}$ [see Fig. 7(c) and (d)], which represents the propagation distance weighted by the instability parameter $\eta_{t}$. The high-$\eta_{t}$ configuration (d) gives rise to higher (roughly by a factor 2) oscillation-amplitude growth compared to the optimal-$\eta_{t}$ configuration (c) at the end of the propagation distance. Note that while the latter sees more energy gain, there is less charge; the overall energy extracted from the wake is the same in both cases. Importantly, both configurations follow the analytical expression in Eq. 3 closely.
@@ VI.1 Uncorrelated energy spread
We can see that initially, with 1% energy spread, the oscillation-amplitude growth is reduced compared to Fig. 7 with no initial energy spread. For the same accelerating field $E_{z}$, the shorter bunches experience a greater reduction in their oscillation-amplitude growth such that they begin to show a large deviation from the expected growth. This is because the phase mixing of the betatron oscillation kicks in before significant oscillation-amplitude growth has had time to develop; the tails of the longer bunches experience much stronger initial wakefields, which causes a much faster growth rate in the oscillation-amplitude. The same behavior can be observed for the bunches with 2% and 4% energy spread but at a larger scale. Additionally for the optimal-$\eta_{t}$ case at an energy spread of 4%, the phase mixing happens at a sufficiently small timescale compared to the instability such that the oscillation-amplitude growth of the bunches is damped to below 1 before the instability grows.
@@ VI.2 Ion motion
If the electron-beam density in a plasma-based accelerator is large enough, the space-charge fields of the beam induce non-negligible motion of the background ions in the plasma [27]. Recently, it has been shown that ion motion can mitigate the hose instability [28]. Ion motion induces a head-to-tail variation in the focusing force, which induces phase mixing of the betatron oscillations of the electron bunch. Hence, the hose instability or the BBU instability can be suppressed by phase mixing breaking up the coherence of the oscillations. The optimal-$\eta_{t}$ and high-$\eta_{t}$ cases are now compared, with ion motion assuming hydrogen ions included. In these simulations, the transverse fields now contain a component induced by the ions, which increases the measured values of $\eta_{t}/\eta_{t,\mathrm{predicted}}$ to 4.5 and 4.6 (from 1.6 and 3.1, respectively). In order to better compare to the no-ion-motion simulations (Fig. 7), we continue to employ the no-ion-motion values of $\eta_{t}$ when calculating $\mu\eta_{t}$ in Fig. 9. The figure shows that the amplitude growth is reduced compared to the predicted growth for most bunch lengths in both cases. As is the case without ion motion, the case with optimal field value generally sees a smaller amplitude growth than the non-optimal case. Here, the amplitude growth quickly reaches and subsequently oscillates around a plateau—a complex behavior caused by the nonlinear focusing fields of the ion-density spike. In the non-optimal case, the decoherence effect of ion motion is not large enough to counteract the rapid growth of the instability, but the amplitude growth does eventually reduce compared to no ion motion. Choosing the optimal accelerating gradient is therefore essential even in cases where ion motion suppresses the BBU instability.
@@ VII Conclusions
The efficiency–instability relation poses a significant challenge for high-efficiency acceleration of high-quality electron bunches. We find that in the ideal case with an optimally loaded (uniform) accelerating field, as well as no energy spread or ion motion, the previously proposed efficiency–instability relation in Eq. 2 does not in fact represent either universal relation or an upper limit, but instead a lower limit on the strength of the instability for a given power-transfer efficiency.

### v2

@@ I Introduction
To be viable for colliders, the plasma-wakefield acceleration (PWFA) scheme must ensure the accelerated electron bunch is stable enough transversely to achieve high luminosity in collisions. In the blowout regime [10], where a driving electron bunch with a larger density than the background plasma expels all the plasma electrons in its path and forms a focusing and accelerating ion column, transverse stability is limited by the coupling between the beam and the electrons traveling in the sheath of the blowout. High driver-to-trailing-bunch power-transfer efficiency is also required in plasma-accelerator facilities [11] for them to operate cost-effectively. However, increasing the driver-to-trailing-bunch power-transfer efficiency gives rise to larger growth rates of transverse instabilities. These instabilities induce an oscillation in the trailing bunch about the plasma channel axis that rapidly increases during acceleration. Such transverse instabilities lead to emittance growth, oscillation-amplitude growth, and, if not mitigated, significant charge loss in PWFAs.
The first type of transverse instability for the blowout regime in PWFAs was discussed by Whittum et al. [12] and found to have a rapid growth rate. The instability is seeded by transverse asymmetries such as tilts in the beam. The offset beam electrons perturb the path of the plasma electrons traveling in the sheath of the plasma wake. The beam oscillates about the plasma channel axis due to the focusing forces from the exposed ions. The perturbation of the plasma wake and the oscillation of the beam slices are coupled and behave as harmonic oscillators. Each longitudinal slice drives the oscillation of the slice positioned right behind it, leading to a resonance behavior whose amplitude increases exponentially towards the tail of the beam. If not mitigated, this growth continues exponentially with the propagation distance. The instability was originally discussed for single beams in plasma (i.e., the drive beam), in which case it is known as the hose instability due to its firehose-like behavior. This instability was later found to be mitigated by the inherent energy spread developed in the driver during propagation [13], as well as the non-linear focusing fields present when the transverse beam size is large [14]. However, these mitigation techniques are not sufficient for the trailing bunch, as future collider applications typically require sub-percent energy spread [11, 15], and the trailing bunch must be matched to small bunch sizes to preserve emittance [16, 8]. When discussing the trailing bunch, this instability is often referred to as the beam-breakup instability and is conceptually similar to that discussed in conventional RF accelerators since the 1960s [17].
@@ II Mapping the parameter space with only three variables
We aim to scan the entire parameter space that affects the strength of the BBU instability. The parameters chosen can be scaled to any plasma density. We turn off ion motion and suppress the energy spread by only working with optimally loaded wakefields and assuming a cold plasma. We transversely offset the trailing bunch by 0.01/$k_{p}$, sufficiently small to perturb the shape of the plasma wake only weakly. The plasma wake has a thin electron sheath [22], such that at the longitudinal position where the blowout is at its largest, the only contribution of the driver is to determine the maximum radius of the blowout $R_{b}$. This is the first parameter we consider in our parameter mapping. The second parameter to consider is the strength of the accelerating field, which is determined by the longitudinal position of the head of the trailing bunch in the plasma wake for a given $R_{b}$. The third parameter is the bunch length of the accelerated bunch. When we scan the bunch length, the charge of the trailing bunch is not a free parameter, as there is only one current profile that can load a perfectly uniform field at a given accelerating gradient and wake radius. The parameters that we include in our grid are, therefore: (1) the normalized wake radius $R_{b}k_{p}$; (2) the normalized accelerating field $E_{z}/E_{0}$; and (3) the bunch length of the trailing bunch.
Figure 1: Plasma-electron density (blue) and beam-electron density (orange) from a particle-in-cell simulation. The solid black line shows the flattened longitudinal wakefield. Different colored lines indicate different bunch lengths that can be compared within the same simulation. For a given loading of the accelerating field, longer bunches will reach higher drive-to-trailing-bunch power-transfer efficiency. The beam is offset by 0.01/$k_{p}$. The total bunch length extends to the end of the blowout.
Figure 2: Grid schematic illustrating the parameter space (selected points only shown). In one dimension, we scan over the blowout wake size $R_{b}k_{p}$. The driver changes only in current along the $R_{b}k_{p}$ direction. In the other dimension, we scan over the normalized accelerating field that the trailing bunch is sampling $E_{z}/E_{0}$. Along the $E_{z}/E_{0}$ direction, the driver is identical, but the longitudinal offset of the trailing bunch changes and the beam current profile is adjusted to beam load optimally. Each simulation also provides a scan over the bunch length. The top left inset in the grey region shows the full range of achievable normalized accelerating fields for each value of $R_{b}k_{p}$ in green.
@@ IV Results
As the value of $E_{z}/E_{0}$ increases and approaches the optimal value, the value of $\eta_{t}/\eta_{t,\mathrm{predicted}}$ decreases at a rate that is roughly constant in log-log space. However, as the value of $E_{z}/E_{0}$ passes the optimal value and begins to increase again, the value of $\eta_{t}$ becomes larger at an increasing rate. In particular, $\eta_{t}/\eta_{t,\mathrm{predicted}}$ for the cases with fields larger than the optimal, increases exponentially when the power-transfer efficiency $\eta_{p}$ tends towards 1. This is shown in Fig. 4 for the normalized maximum blowout radii of $R_{b}k_{p}=0.99$ and $R_{b}k_{p}=2.55$ and power-transfer efficiencies of 5%, 20% and 50%. We fit the $\eta_{t}/\eta_{t,\mathrm{predicted}}$ data points with values of $E_{z}/E_{0}$ less than the optimal value $E_{z,\mathrm{opt}}$, which is obtained when $\eta_{t}/\eta_{t,\mathrm{predicted}}$ is minimized, against $E_{z}/E_{0}$ and $R_{b}k_{p}$ for a generic power-law relationship. The resulting equation
@@
gives very good agreement with the simulated results.
@@
This equation indicates the optimal accelerating field for minimizing transverse instability, and works for efficiencies up to $\sim$60% [^1] We exclude the two highest $R_{b}k_{p}$ values to improve the fit.. The optimal values of $E_{z}/E_{0}$ found from simulation do not deviate by more than 10% at low efficiency and 14% at high efficiency from the predicted values given by Eq. 9.
@@ V Amplitude growth
The oscillation-amplitude growth of the tail of the trailing bunch in each case is evaluated against the normalized quantity $\mu\eta_{t}$ [see Fig. 7(c) and (d)], which represents the propagation distance weighted by the instability parameter $\eta_{t}$. The high-$\eta_{t}$ configuration (d) gives rise to higher (roughly by a factor 2) oscillation-amplitude growth compared to the optimal-$\eta_{t}$ configuration (c) at the end of the propagation distance. Note that while the latter sees more energy gain, there is less charge such that the efficiency, as defined in Eq. 6, is the same for both configurations; the overall energy extracted from the wake is the same in both cases. Importantly, both configurations follow the analytical expression in Eq. 3 closely.
@@ VI.1 Uncorrelated energy spread
We can see that initially, with 1% energy spread, the oscillation-amplitude growth is reduced compared to Fig. 7 with no initial energy spread. For the same accelerating field $E_{z}$, the shorter bunches experience a greater reduction in their oscillation-amplitude growth such that they begin to show a large deviation from the expected growth. This is because the phase mixing of the betatron oscillation kicks in before significant oscillation-amplitude growth has had time to develop; the tails of the longer bunches experience much stronger wakefields, which causes a much faster growth rate in the oscillation-amplitude. The same behavior can be observed for the bunches with 2% and 4% energy spread but at a larger scale. Additionally for the optimal-$\eta_{t}$ case at an energy spread of 4%, the phase mixing happens at a sufficiently small timescale compared to the instability such that the oscillation-amplitude growth of the bunches is damped to below 1 before the instability grows.
@@
As mentioned in the introduction, future collider applications typically require sub-percent energy spread after acceleration. This means that only the beams in Figs. 8(a–d) are close to being viable for a collider application. However, this simulation only considers a single plasma stage. For a collider with several plasma stages, there are additional effects that would have to be taken into account, such as longitudinal movement of the electrons between stages [27]. Additionally, only the energy spread at the end of the acceleration stages is strictly constrained. Therefore, one could possibly have a large energy spread in the beam when going through the early stages of the accelerator, where the oscillation amplitude grows the fastest. Finally, as discussed in the next section, in collider applications the energy spread may be combined with ion motion to increase the damping.
@@ VI.2 Ion motion
If the electron-beam density in a plasma-based accelerator is large enough, the space-charge fields of the beam induce non-negligible motion of the background ions in the plasma [28]. Recently, it has been shown that ion motion can mitigate the hose instability [29]. Ion motion induces a head-to-tail variation in the focusing force, which induces phase mixing of the betatron oscillations of the electron bunch. Hence, the hose instability or the BBU instability can be suppressed by phase mixing breaking up the coherence of the oscillations. The optimal-$\eta_{t}$ and high-$\eta_{t}$ cases are now compared, with ion motion assuming hydrogen ions included. In these simulations, the transverse fields now contain a component induced by the ions, which increases the measured values of $\eta_{t}/\eta_{t,\mathrm{predicted}}$ to 4.5 and 4.6 (from 1.6 and 3.1, respectively). In order to better compare to the no-ion-motion simulations (Fig. 7), we continue to employ the no-ion-motion values of $\eta_{t}$ when calculating $\mu\eta_{t}$ in Fig. 9. The figure shows that the amplitude growth is reduced compared to the predicted growth for most bunch lengths in both cases. As is the case without ion motion, the case with optimal field value generally sees a smaller amplitude growth than the non-optimal case. Here, the amplitude growth quickly reaches and subsequently oscillates around a plateau—a complex behavior caused by the nonlinear focusing fields of the ion-density spike. In the non-optimal case, the decoherence effect of ion motion is not large enough to counteract the rapid growth of the instability, but the amplitude growth does eventually reduce compared to no ion motion. Choosing the optimal accelerating gradient is therefore essential even in cases where ion motion suppresses the BBU instability. For ion motion to be viable as a damping mechanism in collider applications, one must prove that the instability is significantly damped and that emittance is preserved.
@@ VII Conclusions
The efficiency–instability relation poses a significant challenge for high-efficiency acceleration of high-quality electron bunches. We find that in the ideal case with an optimally loaded (uniform) accelerating field, as well as no energy spread or ion motion, the previously proposed efficiency–instability relation in Eq. 2 does not in fact represent either universal relation or an upper limit, but instead a lower limit on the strength of the instability for a given power-transfer efficiency. In short, we update the efficiency–instability relation to instead be an inequality given by
$\eta_{t}\geq\frac{\eta_{p}^{2}}{4(1-\eta_{p})},$ (10)
which holds assuming no energy spread or ion motion.
@@
The initial energy spread needed for significant damping of the instability was shown to be larger than what is typically allowed in collider applications. However, this is for no ion motion in a single plasma stage. In a plasma accelerator used for collider applications one would need several plasma stages, for which the dynamics of the instability is non-trivial. Additional effects would have to be considered, such as longitudinal movement of the beam electrons between stages. Furthermore, one could combine the initial energy spread with additional damping mechanisms such as ion motion. A natural continuation of the work presented in this article would be to compute a similar parametric mapping, but focused on damping the amplitude growth for both single and multiple plasma stages.